OpenStudy (anonymous):

Write a quadratic equation in the variable x having the given number as solutions. Type the equation in standard form,ax^2+bx+c=0 Solution:3, only solution.

jimthompson5910 (jim_thompson5910):

x=3 or x=3 (note: this is a double root) x-3=0 or x-3=0 (x-3)(x-3)=0 (I'm using the zero product property here) x^2-6x+9=0 So the answer is \[\large x^2-6x+9=0\]

OpenStudy (anonymous):

2x^2+x-27=0 can be an answer. there are infinite possible anwers, actually :)

OpenStudy (anonymous):

Or (x - 3)(x - 3) can also be written as \[(x -3)^{2}\] which also gives you \[x ^{2}-6x + 9\]

OpenStudy (anonymous):

I don't see how that hold's true

OpenStudy (anonymous):

Problem states that the only solution is 3

jimthompson5910 (jim_thompson5910):

There are an infinite number of answers, but they are all scalar multiples of \[\large x^2-6x+9\] Unfortunately since \[2x^2+x-27\] is not a scalar multiple of the first expression, this means that this isn't one of them (note: roots of this polynomial involve radicals, so the roots aren't as simple)

OpenStudy (anonymous):

Yes I just checked and 2x^2 + x - 27 is not factorable